The perimeter of a rectangle is 48cm. The length of a rectangle is 3cm longer than twice its width. Find the dimensions of the rectangle.
Answer by math_tutor2020(3835)


Answer: 17 cm by 7 cm

Work Shown

w = width
2w+3 = length

2*(length + width) = perimeter of rectangle
2*(2w+3 + w) = 48
2*(3w+3) = 48
6w+6 = 48
6w = 48-6
6w = 42
w = 42/6
w = 7 cm is the width
2w+3 = 2*7+3 = 17 cm is the length

Check:
perimeter = 2*(length+width) = 2*(17+7) = 2*24 = 48 cm
We have confirmed the answers.